Why is probability useful in biology?
It depends largely on the area of biology that one is talking about, however just about every aspect of biology hinges on some aspect of probability.
In Genetics for example, the probability of random assortment of genes can tell how close two genes are on the chromosome. Likewise, probability of different parental genotypes mating will give an idea of what the f1 generation will look like. Also, if one looks at the probability of random mutation of a single point, one can estimate the conversion of one genotype to another (assuming the point mutation is not lethal and the protein product cannot form correctly).
In other words, it is essential in a variety of ways in Biology.
What would be the probability that two curly haired parents would have a straight haired child?
I don't know the probability for certain really. I do think since it (curly hair) is a recessive gene, then as long as one of the grandparents or great grandparents had it(curly hair), then it would be 50/50. Although I hope someone who knows if curly hair is recessive or dominant will chime in.
My husband has very wavy nearly curly black hair. I have very kinky Celtic copper curls. One daughter has my hair, exactly, but strangely her brows and lashes and body hair are all brown, mine are all translucent blond, except my pubes, which are strawberry blond(TMI?) Her hair is even curlier than mine I think.
Our second daughter has straight sandy blond hair. Dark almost black lashes and brows.
So in our case it was 50/50. Plus some strange surprises, the dark brown lashes, brows, slightly curlier hair, etc in our first child. Sandy blond straight hair, which neither of us have, in our second.
Genetics can be very interesting and tricky. You cannot underestimate the importance of genetic traits that may be recessive or have not appeared in a few generations, popping back up.
What do you mean by probability?
Probability is a measure of the expectation that an event will occur or a statement is true. Probabilities are given a value between 0 (will not occur) and 1 (will occur).[1] The higher the probability of an event, the more certain we are that the event will occur.
In a standard deck of 52 playing cards what is the probability of drawing an ace on one draw?
The probability of drawing one of the four aces on one draw in a standard deck of 52 cards is 1 in 13 (7.7 percent).
This is calculated as such: there are 4 ace cards in a deck, which means the chances are 4/52, which reduces to 1/13.
Because you are replacing the marbles then it is an independent event.
P(1st one is not green) = 1 - P(first green),
equally P(2nd one is not green) = 1 - (second green),
Thus it reads P(¬G ^ ¬G) = P(¬G) * P(¬G)
= 15/20 * 15/20
= 225/400
= 9/16
If there is a 30 chance of rain on tomorrow what is the probability it will not rain on tomorrow?
There is a 70% chance it will not rain tommorrow! There is a 70% chance it will not rain tommorrow!
What is the probability of getting heads on all 3 tosses if a fair coin is tossed 3 times?
1/2 x 1/2 x 1/2 = 1/8 The probability is 1 chance in 8 sequences, or .125
What is the probability of getting 2 tails if a coin is tossed 4 times?
1/2*1/2*1/2*1/2+1/2*1/2*1/2*1/2+1/2*1/2*1/2*1/2+1/2*1/2*1/2*1/2+1/2*1/2*1/2*1/2+1/2*1/2*1/2*1/2+1/2*1/2*1/2*1/2+1/2*1/2*1/2*1/2= 1/2
(8C3 x 9C4)/17C7= (56 x 126)/19448= about 36.3%
Brad flips a coin 3 times what is probability it will land heads up all 3 times?
assuming it's the coin that can land on "heads" and not Brad, also assuming that the coin is perfectly uniformly dense and has a uniform shape and the thing the coin lands on is perfectly still (probably not Brad's head), and has a constant coefficient of friction for wherever the coin touches it, and a constant air friction coefficient, I would say 3/8. or we can accept the fact that there are two sides to a coin, the probability of one side landing up is 1/2, if rolled 3 times, probablility = 3 * .5 *.5 *.5 just perform an experiment
What is the total amount of your purchase if you bought 5 shirts at $m and 2 more at 50% off.
How will probability help me in life?
Probability is everywhere! If you are entering a raffle, wouldn't you like to know what the probability of you winning is? Or perhaps the lottery? If you need to pick a certain color marble out of a bag of marbles to win a prize, wouldn't you like to know the probability of picking the winning color? Probability is everywhere and can help you in life.
it is a fair chance so 1/2 :P
http://mathforum.org/dr.math/faq/faq.prob.world.html
I am unsure of the answer to this question, but I feel the need to remove the previous ridiculous answer stating that "women with down syndrome are sterile." They ARE indeed fertile.
I am in search of the same information, and will be editing this post once I find the correct answer.
most likely
A coin is tossed 3 times. What is the probability of getting 3 tails in a row?
The answer I'm editing says the odds are 1 in 8. This is true only if you actually mean the probability of getting 3 tails in a row, rather than just 3 of either heads or tails in a row.
In mentioned case, the first flip doesn't matter which side it lands on, just the proceeding two flips do.
So, the odds of flipping a coin three times with the same outcome are 1 in 2^2 or 1 in 4.
The odds of flipping three tails in a row are 1 in 2^3 or 1 in 8.
How many combinations can be formed from 12 bits?
For every bit, there are two combinations.
For 12 bits, you need to multiply all together:
2 x 2 x 2 x ... x 2, or 2 to the power 12, which is equal to 4096.
What is the probability of drawing a heart from a normal deck of 52 cards?
A deck of cards consists of four different suits: hearts, clubs, spades and diamonds. Each suit is represented evenly in the deck, so the odds of drawing a heart are 1 in 4.
The probability of guessing any one is 1 out of 4, or 0.25.
Assume that the choices are made independently. Then, if X is the random variable which represents the number of successes (correct guesses), X is a Binomial variable with n = 5 and p = 0.25.
Then Prob(X = 3) = 5C3*p^3*(1-p)^(n-3)
= 10*(0.25)^3*(0.75)^2
= 0.088, approx.
What_is_the_probability_of_spinning_a_free_on_one_spin_and_an_s_on_your_second_if_you_have_a_spinner_with_1_space_labeled_free_and_the_remaining_11_mississippi_with_one_letter_in_each.
So, you would multiply 1/12 times 4/12 to get 4/144 or 1/36
What is the probability of drawing an ace in a deck of 52 cards?
there are 4 aces in a deck, and 52 cards, so the probability is 4/52